OpenStudy (anonymous):

solve the proportions: -3/x=2/8

OpenStudy (kropot72):

First multiply both sides of the equation by x to remove x from the denominator. Can you do this?

OpenStudy (anonymous):

x= -12

OpenStudy (anonymous):

so its -3/1=2/9?

OpenStudy (anonymous):

2/8*

OpenStudy (anonymous):

2/8*

OpenStudy (kropot72):

\[-\frac{3}{x}\times x=-3=\frac{2}{8}\times x=\frac{2x}{8}\] \[-3=\frac{2x}{8}\] Now multiply both sides of this equation by 8: -(3 * 8) = 2x \[x=-\frac{24}{2}=?\]